# MindTuner: Cross-Subject Visual Decoding with Visual Fingerprint and Semantic Correction

### v1

@@ Abstract.
Decoding natural visual scenes from brain activity has flourished, with extensive research in single-subject tasks and, however, less in cross-subject tasks. Reconstructing high-quality images in cross-subject tasks is a challenging problem due to profound individual differences between subjects and the scarcity of data annotation. In this work, we proposed MindTuner for cross-subject visual decoding, which achieves high-quality and rich-semantic reconstructions using only 1 hour of fMRI training data benefiting from the phenomena of visual fingerprint in the human visual system and a novel fMRI-to-text alignment paradigm. Firstly, we pre-train a multi-subject model among 7 subjects and fine-tune it with scarce data on new subjects, where LoRAs with Skip-LoRAs are utilized to learn the visual fingerprint. Then, we take the image modality as the intermediate pivot modality to achieve fMRI-to-text alignment, which achieves impressive fMRI-to-text retrieval performance and corrects fMRI-to-image reconstruction with fine-tuned semantics. The results of both qualitative and quantitative analyses demonstrate that MindTuner surpasses state-of-the-art cross-subject visual decoding models on the Natural Scenes Dataset (NSD), whether using training data of 1 hour or 40 hours.
conference: ; ;
Figure 1. Cross-Subject Visual Decoding and Image Reconstruction. Subjects with sufficient fMRI data are aligned to decode visual stimuli over a shared network. New subjects with scarce visual stimulus are both aligned to the common space, using the shared model, and introduced their visual fingerprint for efficient decoding.
1. Introduction
@@
Do our brains form unified perceptions when we observe similar objects? Do our unique understandings influence these perceptions differently? The human brain exhibits substantial anatomical similarities in terms of functional organization, including shared attributes like memory, functional connectivity, and visual cortex functions (Chen et al., 2017; Fingelkurts et al., 2005; Stringer et al., 2019). However, individual neural biases always exist due to inherent differences (Wang et al., 2020). Understanding both the similarities and gaps in perception has profound implications for the fields of Artificial Intelligence (AI) (Nie et al., 2023; Zhao et al., 2014) and Brain-Computer Interface (BCI) research (Wang et al., 2009). Visual decoding is a straightforward way to understand the brain, where functional magnetic resonance imaging (fMRI) is a widely embraced non-invasive tool used to decode natural visual stimuli, revealing intricate perceptual and semantic details in the cerebral cortex (Linden, 2021). Consequently, fMRI has garnered considerable attention in image retrieval and reconstruction tasks.
@@
Open-source large-scale fMRI datasets, such as the Natural Scenes Dataset (NSD) (Allen et al., 2022), advance deep learning models to shine in fMRI decoding. Pre-trained cross-modality models like CLIP (Radford et al., 2021) and Stable Diffusion (Rombach et al., 2022) offer effective representation space and models for high-quality visual reconstruction. A large body of literature demonstrates the feasibility of training single-subject decoding models to reconstruct high-fidelity images (Lin et al., 2022; Ozcelik and VanRullen, 2023; Scotti et al., 2023; Chen et al., 2023; Takagi and Nishimoto, 2023; Lu et al., 2023). However, single-subject decoding has drawbacks, including the need to train a unique model for each subject, making it challenging to generalize to new subjects and requiring a substantial amount of fMRI data for training. As is widely recognized, acquiring a large amount of fMRI data for each subject is time-consuming, labor-intensive, and impractical in real-world scenarios. Unfortunately, most current research focuses on single-subject visual decoding rather than exploring the challenging commonalities of the human brain. Consequently, there is an urgent need for cross-subject decoding models that can be effectively transferred to new subjects and perform well in the few-shot setting, as depicted in Figure 1.
@@
The key to achieving cross-subject few-shot decoding lies in effectively utilizing extensive prior knowledge from other subjects or additional modalities. On the one hand, one successful strategy for leveraging knowledge from other subjects involves aligning them to a shared space. Ridge regression is commonly employed for this purpose, aligning voxel inputs from different subjects (Scotti et al., 2024; Ferrante et al., 2023). This approach is preferred due to the low signal-to-noise ratio in fMRI data, where complex non-linear models tend to overfit noise. Nonetheless, the process of visual information perception and generating brain activity in each individual incorporates unique components, referred to as the visual fingerprint (Wang et al., 2020) (refer to Section 3 for more detailed analysis). Current linear alignment methods only enable new subjects to conform to the shared components across subjects, neglecting the perception difference derived from their distinctive visual fingerprint and resulting in limited performance.
@@
On the other hand, an additional strategy involves leveraging multimodal data. Previous alignment methods focused solely on the visual modality, specifically aligning fMRI to the penultimate hidden layer of the Visual Transformer (ViT) as a means of adapting to the inputs of Stable Diffusion. However, this alignment approach is susceptible to slight disturbances, which can lead to semantic errors in the generated images. For visual decoding tasks, the textual modality is highly relevant and has been shown to enhance visual decoding semantically (Scotti et al., 2024). However, previous approaches incorporating text have placed excessive emphasis on directly aligning fMRI with detailed textual descriptions to facilitate the reconstruction process (Takagi and Nishimoto, 2023), which intuitively lacks rationality and yields poor performance. Considering that subjects in visual stimulation fMRI experiments do not have a direct interaction with the textual modality and that individual understanding of visual stimuli varies, the relationship between fMRI and text should be considered implicit.
@@
In this paper, we propose MindTuner: a cross-subject visual decoding framework. Inspired by visual fingerprint in Brain Science (Wang et al., 2020), and with the help of Low-Rank Adaptation (LoRA) for large model lightweight fine-tuning. In correspondence to the above two strategies, we first propose the combination of non-linear Skip-LoRAs and LoRAs to learn the visual fingerprint of new subjects, which are injected into the fMRI encoding network to correct visual perception difference. In addition, we design a Pivot mode that uses images as the central modality to bridge fMRI and text. The Pivot helps to correct the reconstructed image with fine-tuned semantics. Our contributions are summarized as follows:
@@
  • MindTuner makes the first attempt to introduce LoRA as a subject-level fine-tuning module in cross-subject decoding and further design non-linear Skip-LoRA. Their combination shows excellent learning capability for subjects’ visual fingerprint.
@@
  • We introduce a novel fMRI-to-text retrieval paradigm with a Pivot using the image modality. The Pivot achieves optimal fMRI-to-text retrieval accuracy and conducts semantic correction with label prompts for fMRI-to-image reconstruction.
@@
2. Related Work
2.1. Single-Subject Visual Decoding from fMRI
The study of single-subject visual decoding using fMRI in the human brain has been a long-standing endeavor. The core goal of visual reconstruction is to generate higher fidelity images using better fMRI encoding networks that better align to the image representation space. For the image representation space, after the initial research focusing on VGG (Horikawa and Kamitani, 2017; Shen et al., 2019; Du et al., 2023), the image representation of CLIP ViT has significantly been favored due to its own richness and the excellent performance of the diffusion model. For voxel encoding networks of fMRI, initial work was devoted to encoding fMRI using ridge regression due to the low signal-to-noise ratio of fMRI (Takagi and Nishimoto, 2023; Lu et al., 2023; Mai and Zhang, 2023; Ozcelik and VanRullen, 2023). MindEye (Scotti et al., 2023) demonstrated that deep MLP and voxel data augmentation with MixCo can achieve better results, while contrastive learning can assist with semantic matching. In the image reconstruction phase, Brain-Diffuser (Ozcelik and VanRullen, 2023) proposes the collaborative generation of two vision levels, high-level and low-level, corresponding to intact and blurred images, respectively. DREAM (Xia et al., 2024) aligns the fMRI to the depth map and spatial palette of the image and uses a controlled generation model T2I Adapter (Mou et al., 2024), to control the color of the image and low-level visualization. For the text modality in the visual reconstruction task, UniBrain (Mai and Zhang, 2023) directly aligns fMRI to the text representation via ridge regression and then completes the brain caption task with the text generation model. MindEye2 (Scotti et al., 2024) obtains a better caption by placing the generated image representation into the image captioning model, which is used to smooth the generated image. However, both methods either do direct fMRI-to-text alignment or direct image-text alignment, ignoring the role of fMRI. It does not model the indirect relationship between fMRI and text from the perspective of intuitive understanding. What’s more, the single-subject models described above are overly dependent on a large amount of fMRI data, ignoring the real-life scenario of an insufficient amount of single-subject data.
2.2. Cross-Subject Functional Alignment
@@
Functional alignment of different brains is considered to be more effective in assisting downstream tasks than anatomical alignment. Previous functional alignment methods were mainly divided into two perspectives: fMRI data itself and downstream tasks. Among a series of methods starting from the fMRI perspective, Bazeille et al. (Bazeille et al., 2019) and Thual et al. (Thual et al., 2023) minimize an optimal transport cost between voxels of different brains. Huang et al. (Huang et al., 2022) learn shared features between subjects by neural manifolds. The methods based on fMRI self-supervision (Chen et al., 2023; Qian et al., 2023) emphasize obtaining common latent representations between subjects through autoencoders. The advantage of the above methods is that they can perform zero-shot learning, but the disadvantage is that they often perform poorly on downstream tasks, such as visual reconstruction tasks. On another cross-subject alignment research path, researchers are committed to aligning fMRI from different subjects to the shared space and using downstream tasks to monitor functional alignment. The advantage of this type of method is that it can achieve good downstream task performance but can only perform few-shot learning. Most of these methods only use linear models to align the inputs of the model, as overly strong non-linear models have been proven to be more prone to overfitting. Linear fitting (Ferrante et al., 2023) was performed on the responses of subjects to common images, achieving high-quality cross-subject visual reconstruction, but requiring subjects to see the same images. Mindeye2 (Scotti et al., 2024) has improved this and achieved very good results, but it still uses a pure linear alignment head. Nonetheless, these methods only consider the linear relationship between subjects and only fine-tune the input header, which is not flexible.
3. Preliminary on visual fingerprint
@@
Visual stimuli are processed by the Human Visual System(HVS), including the retina, iris, cone cells, etc., and transmitted to the brain to form neural signals. Therefore, as stated in previous research (Xia et al., 2024), brain responses such as fMRI are closely linked to our visual system. Although presented in a systematic manner, there are still significant differences in the visual systems of different individuals. Research of visual fingerprint emphasizes that idiosyncratic biases exist in the underlying representation of visual space propagate across varying levels of visual processing (Wang et al., 2020). Using a position-matching task will find stable subject-specific compressions and expansions within local regions throughout the visual field. As shown on the left of Figure 2, in experiments, subjects were fixated at the center, and a target was displayed briefly at one of five possible eccentricities (depicted by dashed lines, which were not visible in the experiment). After the target disappeared, subjects moved the cursor to match the target’s location. Linear models were used to fit the Distortion Indices(DI), which measured the degree of spatial distortion between subjects:
(1) $DI_{self}$ $\sim\beta_{0}+\beta_{1}\times self,$
@@
As shown on the right of Figure 2, the experiment results show that subjects have their own visual fingerprint, which contains both linear and non-linear components. Within-subject similarity (r=0.71) is significantly higher than between-subject similarity (r=0.22), suggesting that each individual subject has their own unique spatial distortions that are consistent within themselves and distinguished from others.
@@
Figure 2. visual fingerprint experiments across subjects. ’Within’ denotes Pearson correlation coefficient of Distortion Indices in within-subject experiments, while ’between’ denotes between-subject.
4. Method
Figure 3. Schematic diagram of MindTuner for new-subject fine-tuning. All shared modules that have undergone multi-subject pre-training are frozen, including MLP Backbone, and submodules of retrieval, high-level, and low-level. The modules that need to be trained include Linear heads, Skip-LoRAs, Adaptive Projector, and LoRA blocks in linear layers. Images are reconstructed using diffusion prior and then SDXL, followed by semantic correction using SDXL Image-Variation.
@@ 4. Method
The task of cross-subject visual decoding consists of two parts: multi-subject pre-training and new-subject fine-tuning. Given a neural dataset with brain activities of subject $s$, it involves the reconstruction of visual images $I\in\mathbb{R}^{3\times H\times W}$ through a core pipeline whose inputs are flattened and aligned fMRI voxels $V\in\mathbb{R}^{1\times d_{s}}$ after ROI extraction, where $d_{s}$ denotes voxel numbers. To simplify notation, each $(I,V)$ denotes data from original subjects for Section 4.1 and from new subjects for Section 4.2. Our goal is to optimize the $\mathcal{F(\cdot)}$, so that $\mathcal{F}(V)=\hat{I}$, where $\hat{I}$ best approximates $I$. New-subject fine-tuning adheres to the same path, albeit with scarce data. For the multi-subject pre-training, we follow the pipeline of MindEye2 (Scotti et al., 2024), while for the new-subject fine-tuning, we plug into visual fingerprint and Pivot as shown in Figure 3.
4.1. Multi-Subject Pre-training
4.1.1. Multi-Subject Functional Alignment
@@
Note that different subjects have different brain structures and that the number of obtained voxels also varies. Therefore, a mapping model is needed to align the voxel inputs from different subjects to the same dimension as inputs to the model in the phase of multi-subject pre-training. Previous methods emphasized zero-shot using optimal transport models or few-shot using linear regression models. Here, we conduct linear function alignment to learn shared-subject fMRI latent space $M\in\mathbb{R}^{1\times d_{0}}$($d_{0}$ denotes shared input dimension) using subject-specific ridge regression as below:
(2) $M=\textbf{Ridge}^{(s)}(V).$
4.1.2. MLP Backbone
@@
To perform high-fidelity pixel-level reconstruction, rich CLIP image embedding is needed. We use OpenCLIP ViT-bigG/14 for alignment whose image token embedding dimension is $256\times 1664$. Mapped inputs $M$ are fed into an MLP backbone with four residual blocks and a tokenization layer going from $d_{0}$-dim to $256\times 1664$. The MLP Backbone $\epsilon(\cdot)$ serves to convert the fMRI to the intermediate backbone embedding space: $Z=\epsilon(M)$. Note that all subjects shared the same MLP backbone after multi-subject functional alignment. Next, the backbone embeddings are passed into three submodules for retrieval, high-level reconstruction and low-level reconstruction.
4.1.3. Retrieval Submodules
@@
To perform the retrieval task, a simple way is conducting a shallow mapping of the backbone embeddings and supervision with fMRI-to-image CLIP contrastive loss. For scarce fMRI samples, appropriate data augmentation helps model convergence. A recently effective voxel mixture paradigm based on MixCo (Kim et al., 2020) works well. Two raw fMRI voxels $V_{i}$ and $V_{j}$ are mixed into $V_{mix_{i,j}}$ using a factor $\lambda$ sampled from the Beta distribution:
(3) $V_{mix_{i,j}}$ $=\lambda_{i}V_{i}+(1-\lambda_{i})V_{j},$
@@
(4) $\mathcal{L}_{MixCo}=-\frac{1}{|B|}\sum_{i=1}^{|B|}[\lambda_{i}\log\frac{\exp(Z_{mix_{i,j}}\cdot\ g_{i}/{\tau})}{\sum_{m=1}^{|B|}\exp(Z_{mix_{i,j}}\cdot\ g_{m}/{\tau})}$
$+(1-\lambda_{i})\log\frac{exp(Z_{mix_{i,j}}\cdot\ g_{j}/{\tau})}{\sum_{m=1}^{|B|}\exp(Z_{mix_{i,j}}\cdot g_{m}/{\tau})}],$
@@
4.1.4. Low-level and High-level Submodules
Low-level pipeline is widely used to enhance low-level visual metrics in reconstruction images, which map voxels to the latent space of Stable Diffusion’s variational autoencoder (VAE) and is a ’guess’ for the reconstruction. It consists of a MLP and a CNN upsampler with L1 loss in Stable Diffusion’s latent space $z$.
(5) $\mathcal{L}_{lowlevel}=\frac{1}{|B|}\sum_{i=1}^{|B|}|z_{i}-\hat{z_{i}}|.$
On the other hand, high-level pipeline emphasizes pixel-wise alignment. Inspired by DALLE·2 (Ramesh et al., 2022), a diffusion prior is recognized as an effective means of transforming backbone embeddings into CLIP ViT image embeddings, in which mean square error loss is used:
(6) $\mathcal{L}_{prior}=\frac{1}{|B|}\sum_{i=1}^{|B|}||g_{i}-\hat{g_{i}}||^{2}_{2}.$
@@ 4.1.4. Low-level and High-level Submodules
(7) $\mathcal{L}_{multi}=\mathcal{L}_{prior}+\alpha_{1}\mathcal{L}_{lowlevel}+\alpha_{2}\mathcal{L}_{BiMixCo}.$
4.2. New-Subject Fine-tuning
4.2.1. Low-Rank Adaptation
Previous works have demonstrated the effectiveness of low-rank adaptation in large language model fine-tuning with significantly fewer trainable parameters. It is well matched to our cross-subject decoding task for two reasons. One is because most of the current mainstream models for decoding fMRI are MLP-based models, which contain a large number of linear layers, while LoRA has also been shown to achieve good fine-tuning results in the linear layers. Second, in cross-subject scenarios, fMRI data from new subjects are usually scarce, and full-tuning the whole model is usually difficult to grasp, leading to a certain degree of overfitting. For each pre-trained weight matrix $\in\mathbb{R}^{d_{in}\times d_{out}}$ in the multi-subject model, where $d_{in}$ denotes input dimension and $d_{out}$ denotes output dimension, gradient update is constrained with a low-rank decomposition for new-subject adapter matrix $\Delta W$:
(8) $W+\Delta W=W+BA,$
@@ 4.2.1. Low-Rank Adaptation
4.2.2. non-linear Skip-LoRAs

### v2

@@ Abstract
Decoding natural visual scenes from brain activity has flourished, with extensive research in single-subject tasks and, however, less in cross-subject tasks. Reconstructing high-quality images in cross-subject tasks is a challenging problem due to profound individual differences between subjects and the scarcity of data annotation. In this work, we proposed MindTuner for cross-subject visual decoding, which achieves high-quality and rich semantic reconstructions using only 1 hour of fMRI training data benefiting from the phenomena of visual fingerprint in the human visual system and a novel fMRI-to-text alignment paradigm. Firstly, we pre-train a multi-subject model among 7 subjects and fine-tune it with scarce data on new subjects, where LoRAs with Skip-LoRAs are utilized to learn the visual fingerprint. Then, we take the image modality as the intermediate pivot modality to achieve fMRI-to-text alignment, which achieves impressive fMRI-to-text retrieval performance and corrects fMRI-to-image reconstruction with fine-tuned semantics. The results of both qualitative and quantitative analyses demonstrate that MindTuner surpasses state-of-the-art cross-subject visual decoding models on the Natural Scenes Dataset (NSD), whether using training data of 1 hour or 40 hours.
Introduction
@@
Do our brains form unified perceptions when we observe similar objects? Do our unique understandings influence these perceptions differently? The human brain exhibits substantial anatomical similarities in terms of functional organization, including shared attributes like memory, functional connectivity, and visual cortex functions (Chen et al. 2017; Fingelkurts, Fingelkurts, and Kähkönen 2005; Stringer et al. 2019). However, individual neural biases always exist due to inherent differences (Wang, Murai, and Whitney 2020). Understanding both the similarities and gaps in perception has profound implications for the fields of Artificial Intelligence (AI) (Nie et al. 2023; Zhao et al. 2014) and Brain-Computer Interface (BCI) research (Wang et al. 2009). Visual decoding is a straightforward way to understand the brain, where functional magnetic resonance imaging (fMRI) is a widely embraced non-invasive tool used to decode natural visual stimuli, revealing intricate perceptual and semantic details in the cerebral cortex (Linden 2021). Consequently, fMRI has garnered considerable attention in image retrieval and reconstruction tasks.
@@
Open-source large-scale fMRI datasets, such as the Natural Scenes Dataset (NSD) (Allen et al. 2022), advance deep learning models to shine in fMRI decoding. Pre-trained cross-modality models like CLIP (Radford et al. 2021) and Stable Diffusion (Rombach et al. 2022) offer effective representation space and models for high-quality visual reconstruction. A large body of literature demonstrates the feasibility of training single-subject decoding models to reconstruct high-fidelity images (Lin, Sprague, and Singh 2022; Scotti et al. 2023; Chen et al. 2023; Takagi and Nishimoto 2023; Lu et al. 2023). However, single-subject decoding has drawbacks, including the need to train a unique model for each subject, making it challenging to generalize to new subjects and requiring a substantial amount of fMRI data for training. As is widely recognized, acquiring a large amount of fMRI data for each subject is time-consuming, labor-intensive, and impractical in practical scenarios. Unfortunately, most current research focuses on single-subject visual decoding rather than exploring the challenging commonalities of the human brain. Consequently, there is an urgent need for cross-subject decoding models that can be effectively transferred to new subjects and perform well in the few-shot setting, as depicted in Figure LABEL:1_hour.
@@
The key to cross-subject few-shot decoding lies in effectively utilizing extensive prior knowledge from other subjects or additional modalities. On the one hand, one successful strategy for leveraging knowledge from other subjects involves aligning them to a shared space. Ridge regression is commonly employed for this purpose, aligning voxel inputs from different subjects (Scotti et al. 2024; Ferrante, Boccato, and Toschi 2023). This approach is preferred due to the low signal-to-noise ratio in fMRI data, where complex non-linear models tend to overfit noise. Nonetheless, the process of visual information perception and generating brain activity in each individual incorporates unique components, referred to as the visual fingerprint (Wang, Murai, and Whitney 2020) (refer to Preliminary for more detailed analysis). Current linear alignment methods only enable new subjects to conform to the shared components across subjects, neglecting the perception difference derived from their distinctive visual fingerprint and resulting in limited performance.
@@
On the other hand, an additional strategy involves leveraging multimodal data. Previous alignment methods focused solely on the visual modality, as a means of adapting to the inputs of Stable Diffusion. However, this alignment approach is susceptible to slight disturbances, leading to semantic errors in the generated images. For visual decoding tasks, the textual modality is highly relevant and is verified effective in enhancing visual decoding semantically (Scotti et al. 2024). However, previous approaches incorporating text have placed excessive emphasis on directly aligning fMRI with detailed textual descriptions to facilitate the reconstruction process (Takagi and Nishimoto 2023), which intuitively lacks rationality and yields poor performance. Considering that subjects in visual stimulation experiments lack direct interaction with the textual modality and that individual understanding of visual stimuli varies, the relationship between fMRI and text should be considered implicit.
@@
In this paper, we propose MindTuner: a cross-subject visual decoding framework. Inspired by visual fingerprint in Brain Science (Wang, Murai, and Whitney 2020), and with the help of Low-Rank Adaptation (LoRA) for large model lightweight fine-tuning. In correspondence to the above two strategies, we first propose the combination of non-linear Skip-LoRAs and LoRAs to learn the visual fingerprint of new subjects, which are injected into the fMRI encoding network to correct visual perception difference. In addition, we design a Pivot module that uses images as the central modality to bridge fMRI and text. The Pivot helps to correct the reconstructed image with fine-tuned semantics. Our contributions are summarized as follows:
@@
  • MindTuner makes the first attempt to introduce LoRA as a subject-level fine-tuning module in cross-subject decoding and further elaborately design non-linear Skip-LoRA. Their combination shows excellent capability to learn subjects’ visual fingerprint.
@@
  • We introduce a novel fMRI-to-text retrieval paradigm with a Pivot using the image modality. The Pivot conducts semantic correction with label prompts to enhance fMRI-to-image reconstruction.
@@
Related Work
Cross-Subject Functional Alignment
@@
In visual decoding, single-subject models have exposed the issue of excessive reliance on the data volume of individual subjects, leading researchers to shift toward cross-subject studies. Functional alignment of different brains is considered to be more effective than anatomical alignment. Previous functional alignment methods were mainly divided into two perspectives: fMRI data itself and downstream tasks. Among a series of methods starting from the fMRI perspective, Bazeille et al. (Bazeille et al. 2019) and Thual et al. (Thual et al. 2023) minimize an optimal transport cost between voxels of different brains. The methods based on fMRI self-supervision (Chen et al. 2023; Qian et al. 2023) emphasize obtaining common latent representations between subjects through autoencoders. Wills Aligner  (Bao et al. 2024) aligns data from different subjects using the fsaverage template. On another cross-subject alignment research path, most of these methods focus on the shared knowledge among subjects and overlook the complex nonlinear relationships between subjects due to concerns about overfitting. Linear fitting  (Ferrante, Boccato, and Toschi 2023) was performed on the responses of subjects to common images, achieving high-quality cross-subject visual reconstruction, but requiring subjects to see the same images. Mindeye2  (Scotti et al. 2024), MindBridge  (Wang et al. 2024) and UMBRAE  (Xia et al. 2024b) improved this and achieved impressive results, but still losing subject-specific features and ignoring the more complex relationships between subjects, which is not flexible.
Text Modality in Visual Decoding
Decoding visual stimuli from fMRI has been a long-standing endeavor, primarily focusing on the image modality  (Takagi and Nishimoto 2023; Lu et al. 2023; Gong et al. 2024b; Ozcelik and VanRullen 2023; Scotti et al. 2023). However, an increasing number of studies have highlighted the role of text. UniBrain (Mai and Zhang 2023) directly aligns fMRI to the text representation via ridge regression and then completes the brain caption task with the text generation model. MindEye2 (Scotti et al. 2024) obtains a better caption by placing the generated image representation into the image captioning model, which is used to smooth the generated image. NeuroClips utilizes the pretrained BLIP-2  (Li et al. 2023) model to achieve text-assisted video reconstruction  (Gong et al. 2024a). However, these methods either do direct fMRI-to-text alignment or direct image-text alignment, ignoring the indirect relationship between fMRI and text from the perspective of intuitive understanding.
Preliminary on visual fingerprint
@@
Visual stimuli are processed by the Human Visual System(HVS). Therefore, as stated in previous research (Xia et al. 2024a), brain responses such as fMRI are closely linked to our visual system. Although presented in a systematic manner, there are still significant differences in the visual systems of different individuals. Research of visual fingerprint emphasizes that idiosyncratic biases exist in the underlying representation of visual space propagate across varying levels of visual processing (Wang, Murai, and Whitney 2020). Using a position-matching task will find stable subject-specific compressions and expansions within local regions throughout the visual field. As shown on the left of Figure 1, in experiments, subjects were fixated at the center, and a target was displayed briefly at one of five possible eccentricities (depicted by dashed lines, which were not visible in the experiment). After the target disappeared, subjects moved the cursor to match the target’s location. Linear models were used to fit the Distortion Indices(DI), which measured the degree of spatial distortion between subjects:
$DI_{self}$ $\sim\beta_{0}+\beta_{1}\times self,$ (1)
@@
As shown on the right of Figure 1, the experiment results show that subjects have their own visual fingerprint, with both linear and non-linear components. Within-subject similarity (r=0.71) is significantly higher than between-subject similarity (r=0.22), suggesting that each individual subject has their own unique spatial distortions that are consistent within themselves and distinguished from others.
@@
Figure 1: Visual fingerprint experiments across subjects. ’Within’ denotes Pearson correlation coefficient of Distortion Indices in within-subject experiments, while ’between’ denotes between-subject.
Method
Figure 2: Schematic diagram of MindTuner. The training process was split into two phases: Multi-subject Pre-training and Cross-subject Fine-tuning, in which the corresponding modules were trained. The predicted embeddings are first obtained through MindTuner, and then the preliminary reconstructed image was obtained by SDXL unCLIP. The final reconstructed image is obtained by text retrieval and semantic correction by SDXL Image-Variation.
@@
The task of cross-subject visual decoding consists of two parts: multi-subject pre-training and new-subject fine-tuning. Given a neural dataset with brain activities of subject $s$, it involves the reconstruction of visual images $I\in\mathbb{R}^{3\times H\times W}$ through a core pipeline whose inputs are flattened and aligned fMRI voxels $V\in\mathbb{R}^{1\times d_{s}}$ after ROI extraction, where $d_{s}$ denotes voxel numbers. To simplify notation, each $(I,V)$ denotes data from the original subjects for this section. Our goal is to optimize the $\mathcal{F(\cdot)}$, so that $\mathcal{F}(V)=\hat{I}$, where $\hat{I}$ best approximates $I$. New-subject fine-tuning adheres to the same path, albeit with scarce data. For the multi-subject pre-training, we follow the pipeline of MindEye2, while for the new-subject fine-tuning, we plug into visual fingerprint and Pivot as shown in Figure 2.
Multi-Subject Pre-training
Multi-Subject Functional Alignment
@@
It should be noted that the brain structures of different subjects vary, as do the number of voxels obtained. Consequently, a mapping model is required to align the voxel inputs from different subjects to the same dimension as the inputs to the model in the phase of multi-subject pre-training. Here, we employ linear function alignment to learn shared-subject fMRI latent space $M\in\mathbb{R}^{1\times d_{0}}$($d_{0}$ denotes shared input dimension). This is achieved through subject-specific ridge regression, as detailed below:
$M=\textbf{Ridge}^{(s)}(V)$ (2)
MLP Backbone
@@
In order to achieve high-fidelity reconstruction, it is necessary to utilize a substantial quantity of CLIP image embedding. The OpenCLIP ViT-bigG/14 space is employed for alignment, with an image embedding dimension of $256\times 1664$. Mapped inputs $M$ are fed into an MLP backbone comprising four residual blocks and a tokenization layer which transforms the input from a dimension of $d_{0}$ to $256\times 1664$. The MLP Backbone $\epsilon(\cdot)$ serves to convert the fMRI to the intermediate backbone embedding space: $Z=\epsilon(M)$. It should be noted that all subjects shared the same MLP backbone following multi-subject functional alignment. Subsequently, the backbone embeddings are conveyed into three submodules for retrieval, high-level reconstruction, and low-level reconstruction.
Retrieval Submodules
@@
A straightforward approach to performing the retrieval task is to conduct a shallow mapping of the backbone embeddings and supervise it with an fMRI-to-image CLIP contrastive loss. In the case of limited fMRI data, the application of appropriate data augmentation techniques can facilitate the convergence of the model. A recently proposed voxel mixture paradigm, based on MixCo  (Kim et al. 2020), has demonstrated effectiveness. Two raw fMRI voxels $V_{i}$ and $V_{j}$ are mixed into $V_{mix_{i,j}}$ using a factor $\lambda$ sampled from the Beta distribution:
$V_{mix_{i,j}}$ $=\lambda_{i}V_{i}+(1-\lambda_{i})V_{j},$ (3)
@@
$\mathcal{L}_{MixCo}=-\frac{1}{|B|}\sum_{i=1}^{|B|}[\lambda_{i}\log\frac{\exp(Z_{mix_{i,j}}\cdot\ g_{i}/{\tau})}{\sum_{m=1}^{|B|}\exp(Z_{mix_{i,j}}\cdot\ g_{m}/{\tau})}$ (4)
$+(1-\lambda_{i})\log\frac{exp(Z_{mix_{i,j}}\cdot\ g_{j}/{\tau})}{\sum_{m=1}^{|B|}\exp(Z_{mix_{i,j}}\cdot g_{m}/{\tau})}]$
@@
Low-level and High-level Submodules
The low-level pipeline is a widely utilized technique for the enhancement of low-level visual metrics in reconstruction images. This involves the mapping of voxels to the latent space of Stable Diffusion’s Variational AutoEncoder (SDVAE), which serves as a surrogate for the reconstruction. The pipeline comprises an MLP and a CNN upsampler with L1 loss in Stable Diffusion’s latent embeddings $z$.
$\mathcal{L}_{lowlevel}=\frac{1}{|B|}\sum_{i=1}^{|B|}|z_{i}-\hat{z_{i}}|$ (5)
Conversely, a high-level pipeline places greater emphasis on semantic alignment. Inspired by DALLE·2 (Ramesh et al. 2022), a diffusion prior is recognized as an effective means of transforming backbone embeddings into CLIP ViT image embeddings, in which mean square error loss is used (further insights on the deployment of diffusion priors can be found in Appendix B.4):
$\mathcal{L}_{prior}=\frac{1}{|B|}\sum_{i=1}^{|B|}||g_{i}-\hat{g_{i}}||^{2}_{2}$ (6)
@@
$\mathcal{L}_{multi}=\mathcal{L}_{prior}+\alpha_{1}\mathcal{L}_{lowlevel}+\alpha_{2}\mathcal{L}_{BiMixCo}$ (7)
Where $\alpha$ denotes the weight to balance multiple losses.
New-Subject Fine-tuning
Low-Rank Adaptation
Previous work has demonstrated the effectiveness of low-rank adaptation in fine-tuning large language models with significantly fewer trainable parameters. It is well suited to our multi-subject decoding task for two reasons. First, most of the current mainstream models for fMRI decoding are MLP-based models that contain a large number of linear layers, whereas LoRA has been shown to achieve good fine-tuning results in the linear layers. Second, in cross-subject scenarios, fMRI data from new subjects are usually scarce, and full-tuning the whole model is usually difficult to grasp, leading to a certain degree of overfitting. For each pre-trained weight matrix $\in\mathbb{R}^{d_{in}\times d_{out}}$ in the multi-subject model, where $d_{in}$ denotes input dimension and $d_{out}$ denotes output dimension, gradient update is constrained with a low-rank decomposition for new-subject adapter matrix $\Delta W$:
$W+\Delta W=W+BA$ (8)
@@
non-linear Skip-LoRAs